OpenStudy (anonymous):

Find F'(X) for F(x)=x^2-x

OpenStudy (ranga):

Take each term and differentiate. Derivative of x^n = n * x^(n - 1)

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

i thought you use the lim as h approaches 0 f(x+h)-f(x)/h

OpenStudy (ranga):

F'(x), y' all mean find the first derivative. y' is same as dy/dx

OpenStudy (anonymous):

yes i know

OpenStudy (ranga):

If you know why are you hesitating taking the derivative?

OpenStudy (math&ing001):

2x-1

OpenStudy (math&ing001):

You guys are complicating the answers

OpenStudy (anonymous):

not really, i know what the answer is but i get stick on one part of it

OpenStudy (ranga):

@math&ing001 You guide people to do the work and find the answers themselves instead of you doing the homework for others. How ill they learn with your method?

OpenStudy (math&ing001):

If they ask about the method you tell them then

OpenStudy (ranga):

What part are you stuck on?

OpenStudy (anonymous):

umm i get stuck on the part where i factor so (x+h)^2-x-(x^2-x)/h

OpenStudy (math&ing001):

[(x+h)^2-x^2]/h=[(x+h-x)*(x+h+x)]=[h*(2x+h)]/h -> 2x

OpenStudy (ranga):

When you are at the stage of calculus where they have only taught you limits and not yet derivatives then you have to do the problem the long way: find the limit as h approaches 0 f(x+h) - f(x) / h. But if you already know derivatives then you can just differentiate each term. It all depends on what stage you are at.

OpenStudy (anonymous):

okay i got the 2x but where does the -1 come from

OpenStudy (math&ing001):

From the second part -x

OpenStudy (anonymous):

so why exactly is it -x?

OpenStudy (math&ing001):

Same method

OpenStudy (math&ing001):

[-(x+h)-(-x)]/h=-h/h=-1

OpenStudy (anonymous):

ohh okay i see now thank you :)

OpenStudy (math&ing001):

Your welcome :)